OpenStudy (anonymous):

HELP PLEASE? The bill at a restaurant is $30. The customer wants to leave a tip of 18%. Part A: Find the total of the bill including the tip. Make sure to show your work in an orderly fashion for credit. Part B: What percent tip would the customer leave if they paid $41.40? Make sure to show all of your work to earn credit

OpenStudy (misssunshinexxoxo):

A. 30 * 1.18 = $35.40 B. (41.40 - 30) / 30 = 11.40 / 30 = .38 = 38%

OpenStudy (misssunshinexxoxo):

The bill inclusive tip is 35.40 $. First you calculate how much are 18 % of 30 $. 30 $ : 100% = 0.3 $ ( = 30 cent) 0.3 $ are 1 % 0.3$ x 18 % = 5.40 $ 30 $ + 5.40 $= 35.40 $ B: 41.40 $ - 30 $= 11.40 $ ( tip) 30 $ = 100 % 0.3 $ = 1 % 11.40 $ : 0.3 = 38 % test : 30 $ ( = 100 % ) + 11.40 $ ( = 38 % ) = 41.40 $ ( 138 % ) 138 % x .03 $ = 41.40 $

OpenStudy (misssunshinexxoxo):

Any questions jbbreezy?